# Quasi-steady emission from repeating fast radio bursts can be explained by magnetar wind nebula

### v1

@@ 5 Observed FRB properties
With the constraints obtained for NS parameters ($P_{i}$, $B_{\rm dip}$, $t_{\rm age}$) in Section 4, we compute the spectral energy distribution (SED) and light curve of the associated PRS, in addition to the time evolution of source DM for the three FRBs to compare them with the radio data. Efficient conversion of rotational energy to particle energy in the termination shock region is required to explain the observed quasi-steady radio emission. We consider two scenarios: (a) rotation-powered model with ($\epsilon_{B}=0.01$, $\gamma_{b}=10^{5}$, $q_{1}=1.5$, $q_{2}=2.5$), and (b) magnetar-flare-powered model with ($\epsilon_{B}=0.1$, $\gamma_{b}=10^{3}$, $q_{1}=q_{2}=2$). We use a opacity $\kappa=0.05\,{\rm cm^{2}/g}$ for the SN ejecta, and set the NS mass $M_{\rm NS}=1.4\,M_{\odot}$, radius $R_{\rm NS}=10\,{\rm km}$ and initial SN explosion radius $R_{0}=10^{6}\,{\rm km}$.
Figure 3: Effect of NS age ($t_{\rm age}$) on the spectral energy distribution of persistent radio emission is shown for FRB 121102 (190520) [201124] in the left (center) [right] panel. Data from radio observations of each PRS source is shown with filled circles. In each panel, we show the results for USSN/CCSN progenitors using solid/dashed curves, for a fixed ($B_{\rm dip}$, $P_{i}$) combination and varying $t_{\rm age}$ – including the best-fit NS age for the respective source. We assume the rotation-powered model with $\epsilon_{B}=0.01$, $\gamma_{b}=10^{5}$, $q_{1}=1.5$ and $q_{2}=2.5$.
Figure 4: The effect of NS dipolar magnetic field $B_{\rm dip}$ (in top row) and initial spin period $P_{i}$ (in bottom row) on the spectral energy distribution of persistent radio emission is shown for FRB 121102 (190520) [201124] in the left (center) [right] column panels. In each panel, the data from radio observations is shown with filled circles and results for USSN/CCSN progenitors are shown using solid/dashed curves. For each FRB, we fix $t_{\rm age}$ to the best-fit value obtained from Figure 3 and vary $B_{\rm dip}$ or $P_{i}$ – including their best-fit combinations for the respective source. As in Figure 3, we assume microphysical parameters corresponding to the rotation-powered model.
@@ 5.1 SED & light curve of the PRS
We first numerically compute the radio SEDs for the nebular emission using the code developed by Murase et al. (2015), which takes into account the effect of electron-positron pairs relevant for emission at $t\gtrsim t_{\rm sd}$. We incorporate the inverse Compton and synchrotron emission processes to solve the time-dependent kinetic equation (3.2), and account for electromagnetic cascades. The external radiation fields are taken to be cosmic microwave background (CMB) and extragalactic background light (EBL). In our calculations, we have included the effect of synchrotron self-absorption (SSA) in magnetised nebula and free-free absorption in the SN ejecta, along with the Razin effect.
@@
Figure 5: Light curves for the persistent radio emission associated with FRB 121102 (190520) [201124] are shown in the left (center) [right] panel. The corresponding data at various radio frequencies for each source is shown using unfilled circles. As earlier, the results for USSN and CCSN progenitors are shown with solid and dashed curves, respectively. We fix the NS parameters $B_{\rm dip}$, $P_{i}$ and $t_{\rm age}$ to their best-fit values obtained from Figures 3 and 4 for the rotation-powered model.
@@
For the rotation-powered model and both progenitor types, the effect of $t_{\rm age}$ on the radio spectral energy distribution (SED) is shown for R1/R2/R3 in the left/center/right panels of Figure 3, with the corresponding radio data shown with filled circles. As expected, the synchrotron flux is suppressed for larger NS age ($t_{\rm age}\gg t_{\rm sd}$) due to considerable adiabatic energy losses at late times. Based on comparison with the observed radio flux, we find a best-fit $t_{\rm age}\approx 20\,(20)\,[8]\,{\rm yr}$ for USSN and $t_{\rm age}\approx 32\,(25)\,[10]\,{\rm yr}$ for CCSN models, corresponding to R1 (R2) [R3]. The low-energy synchrotron flux is highly suppressed due to SSA, especially for the large $M_{\rm ej}$ in case of CCSN progenitors. While the USSN model seems marginally more promising for both R1 and R2 to explain the observed PRS emission at smaller energies ($\nu\sim 1\,{\rm GHz}$), only CCSN model satisfactorily explains the observed emission for R3 due to the large suppression of radio flux at smaller energies.
@@
Next we use the best-fit $t_{\rm age}$ obtained from Figure 3 for each FRB to study the effect of NS parameters on the synchrotron radio flux. In particular, we analyze the effect of dipolar magnetic field $B_{\rm dip}$ on the radio SED as shown in the top-row panels of Figure 4. As in Figure 3, the results for R1/R2/R3 are shown in the left/center/right column panels, with the same representation for progenitor types and radio data shown as filled circles. We find that dipolar field $B_{\rm dip}=3\times 10^{12}\,(5\times 10^{12})\,[5.5\times 10^{13}]\,{\rm G}$ explains the observed radio flux for R1 (R2) [R3]. With further increase in $B_{\rm dip}$, the NS spindown timescale ($t_{\rm sd}\propto B_{\rm dip}^{-2}P_{i}^{2}$) becomes smaller for a given $P_{i}$, which results in energy injection into the nebula at early times. This is then accompanied by significant adiabatic losses which leads to a reduced flux.
@@
The effect of NS spin period $P_{i}$ on the radio SED is shown in the bottom-row panels of Figure 4. As expected, the observed radio flux tends to be larger for a smaller $P_{i}$ as more rotational energy can be extracted from the NS. However, this difference is less noticeable for larger $B_{\rm dip}$ (like in case of R3), as $t_{\rm age}\gg t_{\rm sd}$ results in significant adiabatic energy loss. We find that $P_{i}=3\,(1.5)\,[10]\,{\rm ms}$ best explains the observed radio flux for R1 (R2) [R3]. The low-energy radio flux is suppressed for both fixed $P_{\rm i}$ and fixed $B_{\rm dip}$ cases for the CCSN model, especially for R3, due to stronger SSA attenuation in the nebula. In summary, we find that a NS with $t_{\rm age}\approx 20\,{\rm yr}$, $B_{\rm dip}\approx(3-5)\times 10^{12}\,{\rm G}$ and $P_{i}\approx 1.5-3\,{\rm ms}$ in a USSN progenitor can explain the observed PRS flux for R1 and R2. However, in case of R3, the source is more likely to be a younger NS with $t_{\rm age}\approx 10\,{\rm yr}$, $B_{\rm dip}\approx 5.5\times 10^{13}\,{\rm G}$ and $P_{i}\approx 10\,{\rm ms}$ in a CCSN progenitor. The properties of the central NS and its surrounding ejecta material are consistent with those deduced from previous studies (see e.g., Margalit & Metzger 2018; Zhao & Wang 2021).
@@
Figure 6: Time evolution of ${\rm DM}_{\rm host+ns}$ is shown for FRB 121102 (190520) [201124] in the left (center) [right] panel, by considering contribution from the expanding SN ejecta and MWN. Results for USSN and CCSN models are shown with solid and dashed curves, respectively, and the best-fit NS parameters ($B_{\rm dip}$, $P_{i}$) are the same as those in Figure 5. The shaded yellow region in each panel denotes the host galaxy DM contribution including its uncertainty. For FRB 190520, we also show the previously estimated ${\rm DM}_{\rm host+ns}$, which likely includes contribution from foreground galaxy groups and clusters (Lee et al., 2023). The dashed vertical lines (for the CCSN cases) correspond to minimum $t_{\rm age}$ obtained by imposing $\tau_{\rm sa}<1$ in the MWN and $\tau_{\rm ff}<1$ in the SN ejecta, respectively, with the grey shaded areas signifying the regions that are disallowed by the detectability of PRS.
@@
We next compute the radio light curves associated with synchrotron emission from the magnetized nebula. To study the effect of progenitor model (USSN/CCSN), we use the best-fit $B_{\rm dip}$ and $P_{i}$ values obtained for each FRB source from Figure 4. The light curve data at various frequencies is obtained from: Marcote et al. (2017); Plavin et al. (2022); Rhodes et al. (2023) for R1, Niu et al. (2022); Zhang et al. (2023) for R2, and Bruni et al. (2023) for R3. The corresponding NS age is set to $t_{\rm age}=20\,(20)\,[10]\,{\rm yr}$ from Figure 3. We find that the radio flux predictions from the USSN and CCSN models, for the best-fit NS parameters, are effectively indistinguishable based on current observations for both R1 and R2. However, in case of R3, recent observations at $\nu=6\,{\rm GHz}$ clearly favor the CCSN progenitor model, which also agrees with our results based on the radio SEDs shown in Figure 4. As expected, the radio light curve peaks at a later time and with a smaller synchrotron peak for larger ejecta mass $M_{\rm ej}$ and/or larger SN explosion energy $\mathcal{E}_{\rm sn}$ for the CCSN model. Although the synchrotron flux can reduce with an increase in $M_{\rm ej}$, especially for smaller $t_{\rm age}\sim 10\,{\rm yr}$ as in the case of R3, the spectrum is generally softer due to a smaller synchrotron peak. Lastly, the variation in synchrotron peak flux is marginally higher for CCSN progenitors with a larger SN explosion energy $\mathcal{E}_{\rm sn}$.
@@ 5.2 NS age constraint from DM evolution
where ${\rm DM}_{\rm MW/halo}$ is the contribution from the Milky Way interstellar medium (ISM)/halo, ${\rm DM}_{\rm IGM}$ is the contribution from the intergalactic medium (IGM) and ${\rm DM}_{\rm host}$ is the contribution from the FRB host galaxy including its halo. Here ${\rm DM}_{\rm ns}$ only includes contributions from the MWN and SN ejecta that are near source. While the long-term DM variation is due to the expanding SNR (Yang & Zhang, 2017; Piro & Gaensler, 2018), random fluctuations in the DM can be caused by turbulent motions of filament (Katz, 2021).
From theoretical and data-driven estimates of ${\rm DM}_{\rm MW/halo}$ and ${\rm DM}_{\rm IGM}$, assuming models for the average electron number density in these media (see e.g., Cordes & Lazio 2002, 2003; Planck Collaboration et al. 2016; Prochaska & Zheng 2019; Platts et al. 2020), the inferred ${\rm DM}_{\rm host}+{\rm DM}_{\rm ns}$ in equation (16) indicates significant contributions from the host galaxy disk, circumgalactic medium and near-source medium for most FRBs that have an associated PRS (see e.g., Law et al. 2022). In fact, from previous studies (Chatterjee et al., 2017; Niu et al., 2022; Bruni et al., 2023; Lee et al., 2023), we know that the host galaxy and near-source contribution ${\rm DM}_{\rm host}+{\rm DM}_{\rm ns}\sim 140\,(430)\,[185]\,{\rm pc\,cm^{-3}}$ for R1 (R2) [R3] can be considerable. Using Illustris simulations to model the electron number density along various lines of sight, Zhang et al. (2020) showed that the typical DM contribution from host galaxies at $z\sim 0.2$, that are similar to the three FRBs considered here, can be $\sim 50-250\,{\rm pc\,cm^{-3}}$.
@@
  FRB 121102 USSN ($10^{50}\,{\rm erg}$, $0.1\,M_{\odot}$) $20\,{\rm yr}$ $3\times 10^{12}\,{\rm G}$ $3.0\,{\rm ms}$
@@
FRB 190520 USSN ($10^{50}\,{\rm erg}$, $0.1\,M_{\odot}$) $20\,{\rm yr}$ $5\times 10^{12}\,{\rm G}$ $1.5\,{\rm ms}$
@@
FRB 201124 USSN ($10^{50}\,{\rm erg}$, $0.1\,M_{\odot}$) $8\,{\rm yr}$ $5.5\times 10^{13}\,{\rm G}$ $10\,{\rm ms}$
@@
For a young NS, the magnetised nebula and SN ejecta can be dense enough to provide a large DM which is inconsistent with the observations. DM contribution from the near-source medium should decrease with NS age as the MWN and SN ejecta expand over time. In contrast to the nearly unchangeable DM of FRB 121102 (Hessels et al., 2019; Oostrum et al., 2020; Li et al., 2021), the DM of FRB 190520 decreases with the rate of $\sim 0.1\,{\rm pc\,cm^{-3}}$ per day (Niu et al., 2022). We use detailed calculations of ionization and radiation transport for young rapidly rotating NS to predict contributions from the MWN and SN ejecta to FRB DM (see Kashiyama & Murase 2017; Margalit & Metzger 2018).
@@
Figure 6 shows the time evolution of the inferred FRB host galaxy and near source DM contribution, ${\rm DM}_{\rm host}+{\rm DM}_{\rm ns}$, for R1 (R2) [R3] in the left (center) [right] panel. To analyze the DM contributions from the MWN and expanding SN ejecta, for a given progenitor model i.e. USSN/CCSN, we fix the NS parameters to their best-fit values as in Figure 5. The host galaxy DM contribution (including its relative uncertainty) is shown in each panel with the shaded yellow region; the shaded magenta region for R2 denotes the possibly overestimated ${\rm DM}_{\rm host}$ due to significant contributions from the foreground galaxies and clusters (see e.g, Lee et al. 2023). As expected, the DM contribution from the SN ejecta and magnetized nebula is substantially higher for CCSN progenitor models as they have a larger $M_{\rm ej}$. The contribution from the magnetised nebula to ${\rm DM}_{\rm ns}$ is almost negligible compared to that from the denser SN ejecta.
@@
A lower limit on the source age $t_{\rm age}$ follows from the requirement that the SN ejecta should not be significantly dense to overproduce DM relative to ${\rm DM}_{\rm obs}$. We find the minimum NS age allowed by the near-source DM constraint to be $t_{\rm age,min}\sim 1-3\,{\rm yr}$. The dashed vertical lines and the associated grey-shaded regions in Figure 6 correspond to the condition that the SN ejecta and the magnetized nebula should allow for the propagation of radio waves outside the source environment, and therefore be transparent to free-free absorption ($\tau_{\rm ff}\lesssim 1$) and synchrotron self-absorption ($\tau_{\rm sa}\lesssim 1$), respectively. This condition imposes a somewhat stronger constraint on the minimum NS age with $t_{\rm age,min}\sim 10\,{\rm yr}$ for R1, $\sim 8\,{\rm yr}$ for R2 and $\sim 6.5\,{\rm yr}$ for R3. Our inferred best-fit $t_{\rm age}$ from Figure 3 exceeds this minimum NS age required by the absorption criterion.
5.3 Rotation and magnetic powered models

### v2

@@ 5 Observed FRB properties
With the constraints obtained for NS parameters ($P_{i}$, $B_{\rm dip}$, $t_{\rm age}$) in Section 4, we compute the spectral energy distribution (SED) and light curve of the associated PRS, in addition to the time evolution of source DM for the three FRBs to compare them with the radio data. Efficient conversion of rotational energy to particle energy in the termination shock region is required to explain the observed quasi-steady radio emission. We consider two scenarios: (a) rotation-powered model with ($\epsilon_{B}=0.01$, $\gamma_{b}=10^{5}$, $q_{1}=1.5$, $q_{2}=2.5$), and (b) magnetar-flare-powered model with ($\epsilon_{B}=0.1$, $\gamma_{b}=10^{3}$, $q_{1}=q_{2}=2$). We use an opacity of $\kappa=0.05\,{\rm cm^{2}/g}$ for the SN ejecta, and set the NS mass $M_{\rm NS}=1.4\,M_{\odot}$, radius $R_{\rm NS}=10\,{\rm km}$ and initial SN explosion radius $R_{0}=10^{6}\,{\rm km}$.
Figure 3: Effect of NS age ($t_{\rm age}$) on the spectral energy distribution of PRS emission is shown for FRB 20121102A (20190520B) [20201124A] in the left (center) [right] panel. Data from radio observations of each source is shown with filled circles. In each panel, we show the results for USSN/CCSN progenitors using solid/dashed curves, for a fixed ($B_{\rm dip}$, $P_{i}$) combination and varying $t_{\rm age}$ – including the best-fit NS age for the respective source. We assume the rotation-powered model with $\epsilon_{B}=0.01$, $\gamma_{b}=10^{5}$, $q_{1}=1.5$ and $q_{2}=2.5$.
Figure 4: Effect of NS dipolar magnetic field $B_{\rm dip}$ (in top row) and initial spin period $P_{i}$ (in bottom row) on the spectral energy distribution of PRS emission is shown for FRB 20121102A (20190520B) [20201124A] in the left (center) [right] column panels. In each panel, the data from radio observations is shown with filled circles and results for USSN/CCSN progenitors are shown using solid/dashed curves. For each FRB, we fix $t_{\rm age}$ to the best-fit value obtained from Figure 3 and vary $B_{\rm dip}$ or $P_{i}$ – including their best-fit combinations for the respective source. As in Figure 3, we assume microphysical parameters corresponding to the rotation-powered model.
@@ 5.1 SED & light curve of the PRS
We first numerically compute the radio SEDs for the nebular emission using the code developed by Murase et al. (2015, 2021), which takes into account the effect of electron-positron pairs relevant for emission at $t\gtrsim t_{\rm sd}$. We incorporate the inverse Compton and synchrotron emission processes to solve the time-dependent kinetic equation (3.2), and account for electromagnetic cascades. The external radiation fields are taken to be cosmic microwave background (CMB) and extragalactic background light (EBL). In our calculations, we have included the effect of synchrotron self-absorption (SSA) in magnetised nebula and free-free absorption in the SN ejecta, along with the Razin effect.
@@
Figure 5: Light curves for the persistent radio emission associated with FRB 20121102A (20190520B) [20201124A] are shown in the left (center) [right] panel. The corresponding data at various radio frequencies for each source is shown using unfilled circles. As earlier, the results for USSN and CCSN progenitors are shown with solid and dashed curves, respectively. We fix the NS parameters $B_{\rm dip}$, $P_{i}$ and $t_{\rm age}$ to their best-fit values obtained from Figures 3 and 4 for the rotation-powered model.
@@
For the rotation-powered model and both progenitor types, the effect of $t_{\rm age}$ on the radio spectral energy distribution (SED) is shown for FRB 20121102A/20190520B/20201124A in the left/center/right panels of Figure 3, with the corresponding radio data shown with filled circles. As expected, the synchrotron flux is suppressed for larger NS age ($t_{\rm age}\gg t_{\rm sd}$) due to considerable adiabatic energy losses at late times. Based on comparison with the observed radio flux, we find a best-fit $t_{\rm age}\approx 20\,(20)\,[8]\,{\rm yr}$ for USSN and $t_{\rm age}\approx 32\,(25)\,[10]\,{\rm yr}$ for CCSN models, corresponding to FRB 20121102A (20190520B) [20201124A]. The low-energy synchrotron flux is highly suppressed due to SSA, especially for the large $M_{\rm ej}$ in case of CCSN progenitors. While the USSN model seems marginally more promising for both FRBs 20121102A and 20190520B to explain the observed PRS emission at smaller energies ($\nu\sim 1\,{\rm GHz}$), only CCSN model satisfactorily explains the observed emission for FRB 20201124A due to the large suppression of radio flux at smaller energies.
@@
Next we use the best-fit $t_{\rm age}$ obtained from Figure 3 for each FRB to study the effect of NS parameters on the synchrotron radio flux. In particular, we analyze the effect of dipolar magnetic field $B_{\rm dip}$ on the radio SED as shown in the top-row panels of Figure 4. As in Figure 3, the results for FRB 20121102A/20190520B/20201124A are shown in the left/center/right column panels, with the same representation for progenitor types and radio data shown as filled circles. We find that dipolar field $B_{\rm dip}=3\times 10^{12}\,(5\times 10^{12})\,[5.5\times 10^{13}]\,{\rm G}$ explains the observed radio flux for FRB 20121102A (20190520B) [20201124A]. With further increase in $B_{\rm dip}$, the NS spindown timescale ($t_{\rm sd}\propto B_{\rm dip}^{-2}P_{i}^{2}$) becomes smaller for a given $P_{i}$, which results in energy injection into the nebula at early times. This is then accompanied by significant adiabatic losses which leads to a reduced flux.
@@
The effect of NS spin period $P_{i}$ on the radio SED is shown in the bottom-row panels of Figure 4. As expected, the observed radio flux tends to be larger for a smaller $P_{i}$ as more rotational energy can be extracted from the NS. However, this difference is less noticeable for larger $B_{\rm dip}$ (like in case of FRB 20201124A), as $t_{\rm age}\gg t_{\rm sd}$ results in significant adiabatic energy loss. We find that $P_{i}=3\,(1.5)\,[10]\,{\rm ms}$ best explains the observed radio flux for FRB 20121102A (20190520B) [20201124A]. The low-energy radio flux is suppressed for both fixed $P_{\rm i}$ and fixed $B_{\rm dip}$ cases for the CCSN model, especially for FRB 20201124A, due to stronger SSA attenuation in the nebula. In summary, we find that a NS with $t_{\rm age}\approx 20\,{\rm yr}$, $B_{\rm dip}\approx(3-5)\times 10^{12}\,{\rm G}$ and $P_{i}\approx 1.5-3\,{\rm ms}$ in a USSN progenitor can explain the observed PRS flux for FRBs 20121102A and 20190520B. However, in case of FRB 20201124A, the source is more likely to be a younger NS with $t_{\rm age}\approx 10\,{\rm yr}$, $B_{\rm dip}\approx 5.5\times 10^{13}\,{\rm G}$ and $P_{i}\approx 10\,{\rm ms}$ in a CCSN progenitor. The properties of the central NS and its surrounding ejecta material are consistent with those deduced from previous studies (see e.g., Margalit and Metzger 2018; Zhao and Wang 2021).
@@
Figure 6: Time evolution of ${\rm DM}_{\rm host+ns}$ is shown for FRB 20121102A (20190520B) [20201124A] in the left (center) [right] panel, by considering contribution from the expanding SN ejecta and MWN. Results for USSN and CCSN models are shown with solid and dashed curves, respectively, and the best-fit NS parameters ($B_{\rm dip}$, $P_{i}$) are the same as those in Figure 5. The shaded yellow region in each panel denotes the host galaxy DM contribution including its uncertainty. For FRB 20190520B, we also show the previously estimated ${\rm DM}_{\rm host+ns}$, which likely includes contribution from foreground galaxy groups and clusters (Lee et al., 2023). The dashed vertical lines (for the CCSN cases) correspond to minimum $t_{\rm age}$ obtained by imposing $\tau_{\rm sa}<1$ in the MWN and $\tau_{\rm ff}<1$ in the SN ejecta, respectively, with the grey shaded areas signifying the regions that are disallowed by the detectability of PRS.
@@
We next compute the radio light curves associated with synchrotron emission from the magnetized nebula. To study the effect of progenitor model (USSN/CCSN), we use the best-fit $B_{\rm dip}$ and $P_{i}$ values obtained for each FRB source from Figure 4. The light curve data at various frequencies is obtained from: Marcote et al. (2017); Plavin et al. (2022); Rhodes et al. (2023) for FRB 20121102A, Niu et al. (2022); Zhang et al. (2023) for FRB 20190520B, and Bruni et al. (2023) for FRB 20201124A. The corresponding NS age is set to $t_{\rm age}=20\,(20)\,[10]\,{\rm yr}$ from Figure 3. We find that the radio flux predictions from the USSN and CCSN models, for the best-fit NS parameters, are effectively indistinguishable based on current observations for both FRBs 20121102A and 20190520B. However, in case of FRB 20201124A, recent observations at $\nu=6\,{\rm GHz}$ clearly favor the CCSN progenitor model, which also agrees with our results based on the radio SEDs shown in Figure 4. As expected, the radio light curve peaks at a later time and with a smaller synchrotron peak for larger ejecta mass $M_{\rm ej}$ and/or larger SN explosion energy $\mathcal{E}_{\rm sn}$ for the CCSN model. Although the synchrotron flux can reduce with an increase in $M_{\rm ej}$, especially for smaller $t_{\rm age}\sim 10\,{\rm yr}$ as in the case of FRB 20201124A, the spectrum is generally softer due to a smaller synchrotron peak. Lastly, the variation in synchrotron peak flux is marginally higher for CCSN progenitors with a larger SN explosion energy $\mathcal{E}_{\rm sn}$.
@@ 5.2 NS age constraint from DM evolution
where ${\rm DM}_{\rm MW/halo}$ is the contribution from the Milky Way interstellar medium (ISM)/halo, ${\rm DM}_{\rm IGM}$ is the contribution from the intergalactic medium (IGM) and ${\rm DM}_{\rm host}$ is the contribution from the FRB host galaxy including its halo. Here ${\rm DM}_{\rm ns}$ only includes contributions from the MWN and SN ejecta that are near source. While the long-term DM variation is due to the expanding SNR (Yang and Zhang, 2017; Piro and Gaensler, 2018), random fluctuations in the DM can be caused by turbulent motions of filament (Katz, 2021).
From theoretical and data-driven estimates of ${\rm DM}_{\rm MW/halo}$ and ${\rm DM}_{\rm IGM}$, assuming models for the average electron number density in these media (see e.g., Cordes and Lazio 2002, 2003; Planck Collaboration et al. 2016; Prochaska and Zheng 2019; Platts et al. 2020), the inferred ${\rm DM}_{\rm host}+{\rm DM}_{\rm ns}$ in equation (16) indicates significant contributions from the host galaxy disk, circumgalactic medium and near-source medium for most FRBs that have an associated PRS (see e.g., Law et al. 2022). In fact, from previous studies (Chatterjee et al., 2017; Niu et al., 2022; Bruni et al., 2023; Lee et al., 2023), we know that the host galaxy and near-source contribution ${\rm DM}_{\rm host}+{\rm DM}_{\rm ns}\sim 140\,(430)\,[185]\,{\rm pc\,cm^{-3}}$ for FRB 20121102A (20190520B) [20201124A] can be considerable. Using Illustris simulations to model the electron number density along various lines of sight, Zhang et al. (2020) showed that the typical DM contribution from host galaxies at $z\sim 0.2$, that are similar to the three FRBs considered here, can be $\sim 50-250\,{\rm pc\,cm^{-3}}$.
@@
  FRB 20121102A USSN ($10^{50}\,{\rm erg}$, $0.1\,M_{\odot}$) $20\,{\rm yr}$ $3\times 10^{12}\,{\rm G}$ $3.0\,{\rm ms}$
@@
FRB 20190520B USSN ($10^{50}\,{\rm erg}$, $0.1\,M_{\odot}$) $20\,{\rm yr}$ $5\times 10^{12}\,{\rm G}$ $1.5\,{\rm ms}$
@@
FRB 20201124A USSN ($10^{50}\,{\rm erg}$, $0.1\,M_{\odot}$) $8\,{\rm yr}$ $5.5\times 10^{13}\,{\rm G}$ $10\,{\rm ms}$
@@
For a young NS, the magnetised nebula and SN ejecta can be dense enough to provide a large DM inconsistent with observations. DM contribution from the near-source medium should decrease with NS age as the MWN and SN ejecta expand over time. In contrast to the nearly unchangeable DM of FRB 20121102A (Hessels et al., 2019; Oostrum et al., 2020; Li et al., 2021), the DM of FRB 20190520B decreases with the rate of $\sim 0.1\,{\rm pc\,cm^{-3}}$ per day (Niu et al., 2022). We use detailed calculations of ionization and radiation transport for young rapidly rotating NS to predict contributions from the MWN and SN ejecta to FRB DM (see Kashiyama and Murase 2017; Margalit and Metzger 2018).
@@
Figure 6 shows the time evolution of the inferred FRB host galaxy and near source DM contribution, ${\rm DM}_{\rm host}+{\rm DM}_{\rm ns}$, for FRB 20121102A (20190520B) [20201124A] in the left (center) [right] panel. To analyze the DM contributions from the MWN and expanding SN ejecta, for a given progenitor model i.e. USSN/CCSN, we fix the NS parameters to their best-fit values as in Figure 5. The host galaxy DM contribution (including its relative uncertainty) is shown in each panel with the shaded yellow region; the shaded magenta region for 20190520B denotes the possibly overestimated ${\rm DM}_{\rm host}$ due to significant contributions from the foreground galaxies and clusters (see e.g, Lee et al. 2023). As expected, the DM contribution from the SN ejecta and magnetized nebula is substantially higher for CCSN progenitor models as they have a larger $M_{\rm ej}$. The contribution from the magnetised nebula to ${\rm DM}_{\rm ns}$ is almost negligible compared to that from the denser SN ejecta.
@@
A lower limit on source age $t_{\rm age}$ follows from the requirement that SN ejecta should not be significantly dense to overproduce DM relative to ${\rm DM}_{\rm obs}$. We find the minimum NS age allowed by the near-source DM constraint to be $t_{\rm age,min}\sim 1-3\,{\rm yr}$. The dashed vertical lines and the associated grey-shaded regions in Figure 6 correspond to the condition that the SN ejecta and magnetized nebula should allow for the propagation of radio waves outside the source environment, and therefore be transparent to free-free absorption ($\tau_{\rm ff}\lesssim 1$) and synchrotron self-absorption ($\tau_{\rm sa}\lesssim 1$), respectively. This condition imposes a stronger constraint on minimum NS age with $t_{\rm age,min}\sim 10\,{\rm yr}$ for FRB 20121102A, $\sim 8\,{\rm yr}$ for FRB 20190520B and $\sim 6.5\,{\rm yr}$ for FRB 20201124A. Our inferred best-fit $t_{\rm age}$ from Figure 3 exceeds this minimum NS age required by the absorption criterion.
5.3 Rotation powered vs magnetic powered models